T1⁄4 can be taken as time taken for the concentration of a reactant to drop to 3/4 of its initial value . If rate constant for a first order reaction is k, then t1/4 can be written as

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
From the first order reaction,
Kt = ln (An/Ao) 
In this way, for t1/4 
K.t1/4= ln (4/3) 
= O.287 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)